OpenStudy (anonymous):

Solve the in equality x^2>9

OpenStudy (compassionate):

The only solution is, \[x = \sqrt{9} \]

OpenStudy (anonymous):

and express the solution in terms of intervals

OpenStudy (anonymous):

First, take the square root of both sides. absolute value of {x}>(+/-)3 This is now two separate inequalities... x<-3 and x>3.

OpenStudy (anonymous):

\[\sqrt{x^2}>\sqrt{9}\]\[x>3\]

OpenStudy (anonymous):

The given interval is (-infinity,-3) U (3, infinity)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Make sure when you take the square root of something you remember the absolute value signs and the plus or minus signs! =)

OpenStudy (anonymous):

truth :)

OpenStudy (anonymous):

Oh, yes, thank you

OpenStudy (anonymous):

x > 3. :)